For Al-28, calculate a) the mass defect (use Table 18.4) and b) the binding energy in 

kJ/mol.


1
Expert's answer
2021-06-01T02:46:30-0400

Solution

Al-28 =27.981910 amu

Al has 13 protons

28-13=15

15 neutrons

proton=1.00727647 amu

Neutron=1.00866490 amu

predicted mass =13(1.00727647)+15(1.00866490)

=28.22456761 amu

Mass defect= Predicted mass - Actual mass

28.22457-27.98191= 0.24266 amu


b) Binding energy


E=mc2

1 amu = 1.66054 x 10-27kg

c= 2.99792 x108

(0.24266 x 1.66054 x 10-27) x (2.99792 x 108)

= 3.62149 x10-11 J

convert to KJ/mole


"\frac{3.62149 \times 10^-11}{6.022 \times 10^20}"


=6.01377 x 10-32KJ/mol